OpenStudy (anonymous):

Show that PV=nRT becomes P/p=(R/MW)T where MW=28.97kg/kmol

OpenStudy (mos1635):

\[P V = n R T\] \[P V =(m/MW) R T\] \[P (V/m) = (R/MW) T\] \[P (1/ρ) = (R/Mw) T\] Mw = 28.97 is the wean value if gas is air

OpenStudy (anonymous):

Thank you!